OpenStudy (anonymous):

add 8 1/2, 1 1/3, 8 1/4, 6 1/6

OpenStudy (turingtest):

8 1/2=17/2 1 1/3=4/3 8 1/4=33/4 6 1/6=37/6 now find the lowest common denominator (looks like it should be 12): 17/2+4/3+33/4+37/6=(102+16+99+72)/12=289/12